OpenStudy (anonymous):

Let R be the rectangle with vertices (0,0), (8,0), (0,9), (8,9)?

OpenStudy (anonymous):

http://i.imgur.com/ACQ1WoM.png (screenshot of entire problem)

OpenStudy (anonymous):

i havent done this in a while but i would apply greens theorem to this problem \[\int\limits_{c}^{} F dX = \int\limits_{}^{}\int\limits_{R}^{} <curl F> k dA \] where k is the standard unit vector in the positive z direction

OpenStudy (anonymous):

Can you walk me through it? I have no clue how to do these. Its an online course so I dont get a lot of instruction

OpenStudy (anonymous):

heres a link that might help you i would walk you through this but my knowledge of multi variable calc is a bit fuzzy and i dont want to give you any wrong ideas http://tutorial.math.lamar.edu/Classes/CalcIII/CurlDivergence.aspx